Decide how many solutions this equation has:
x^2 + 3 = 0

Answer by stanbon(75887)

Decide how many solutions this equation has:
x^2 + 3 = 0
---
a = 1 ; b = 0 ; c = 3
----
discriminant = b^2-4ac = 0^2-4*1*3 = -12
-------
Since it is negative there are 2 complex-
number solution.
No Real number solutions.
========================================
Cheers,
Stan H.